OpenStudy (18jonea):

A design on the surface of a balloon is 9 cm wide when the balloon holds 62 cm3 of air. How much air does the balloon hold when the design is 18 cm wide? 496 cm3 124 cm3 407 cm3 248 cm3 @amistre64

OpenStudy (amistre64):

kind of an assumptious question what our scalar?

OpenStudy (amistre64):

9 times what, is 18?

OpenStudy (18jonea):

2

OpenStudy (amistre64):

this is important to the solution, how does the linear scalar effect volume?

OpenStudy (18jonea):

so then you would times 62 by 2

OpenStudy (amistre64):

what is the formula for the volume of a sphere?

OpenStudy (18jonea):

4rsquared pi

OpenStudy (amistre64):

no, volume ....

OpenStudy (amistre64):

area is squared (2 dimensions), volume is cubic (3 dimensions)

OpenStudy (18jonea):

4 3πrcubed

OpenStudy (amistre64):

good, 4/3 pi r^3 r is a linear measure, radius. a scale by 2 gives us 2r Vnew = 4/3 pi (2r)^3 or written another way Vnew = 4/3 pi r^3 (2)^3 Vnew = Vold (2)^3

OpenStudy (18jonea):

so what would i do next

OpenStudy (amistre64):

the calculation of course ....

OpenStudy (amistre64):

how much bigger is the new volume when we have a linear scalar value of 2?

OpenStudy (18jonea):

2

OpenStudy (amistre64):

no, thats not what i wrote at all ....

OpenStudy (18jonea):

8?

OpenStudy (amistre64):

Vold = 4/3 pi r^3 r is a linear measure, radius. a scale by 2 gives us 2r Vnew = 4/3 pi (2r)^3 or written another way Vnew = 4/3 pi r^3 (2)^3 Vnew = Vold (2)^3

OpenStudy (18jonea):

2 x 2 x 2

OpenStudy (amistre64):

yeah, 2^3 = 8

OpenStudy (amistre64):

oh good, it looks like the posts are getting jumbled about .... site might go down soon

OpenStudy (18jonea):

what

OpenStudy (amistre64):

the posts arent in order on my end ....

OpenStudy (amistre64):

old volume * 8 = new stuff

OpenStudy (18jonea):

o i got 496 for the new volume is that right?

OpenStudy (18jonea):

?

OpenStudy (here_to_help15):

Sorry to interupt guys but @amistre64 i need you to check something super super fast someone is testing my knowledge and i dont like it

OpenStudy (18jonea):

@amistre64

OpenStudy (amistre64):

62*8 is correct

OpenStudy (18jonea):

The volumes of two similar solids are 1331 m3 and 343 m3. The surface area of the larger one is 484 m2. What is the surface area of the smaller one? 196 m2 343 m2 1372 m2 none of these @amistre64

OpenStudy (18jonea):

@Here_to_Help15

OpenStudy (18jonea):

ok Thanks @amistre64

OpenStudy (here_to_help15):

Ratio of the volumes is 1331/343 = 3.880 linear dimensions are proportional to the of that, and area is proportional to the square of the linear dimension, so over all, it is an exponent of 2/3 so 3.880^(2/3) = 2.469 so dividing 484 / 2.469, we get 196 m^2

OpenStudy (here_to_help15):

In that space i left open this goes there one sec

OpenStudy (here_to_help15):



OpenStudy (anonymous):

what are we trying to figure out ?

OpenStudy (18jonea):

so what would it be

OpenStudy (18jonea):

@Here_to_Help15

OpenStudy (here_to_help15):

I told you @18jonea look at what i wrote

OpenStudy (18jonea):

o sorry so oits 196m squared

OpenStudy (here_to_help15):

mhmm

OpenStudy (18jonea):

o sorry so oits 196m squared